OpenStudy (anonymous):

the sum of a number 6 times it's reciprocal 5 find numbers

OpenStudy (anonymous):

can you rephrase using understandable english?

OpenStudy (anonymous):

The sum of a number 6 times its reciprocal 5 find the numbers

OpenStudy (anonymous):

..... i'm sorry, but i don't understand that... x + 6/x <-- thats all i can get up too i don't understand that 5

OpenStudy (anonymous):

oh wait you're saying the sum IS 5?

OpenStudy (anonymous):

the sum of a number 6 times its reciprocal is 5

OpenStudy (anonymous):

x + 6/x = 5 (x^2 + 6)/ x = 5 x^2 + 6 = 5x x^2 - 5x + 6 = 0 (x - 2)(x - 3) = 0 x = 2, 3 check: 2 + 6(1/2) = 5 3 + 6(1/3) = 5 yup x = 2, 3